# Novel method for in-situ drift velocity measurement in large volume TPCs: the Geometry Reference Chamber of the NA61/SHINE experiment at CERN

### v1

@@ Appendix B Systematic errors and alignment calibration of the TPCs
which is nothing but the already known Eq.(A.1), re-expressed in the notations of this appendix. Therefore, the remaining corrections $t_{0,\mathrm{corr}}$ and $y_{0}$ can be determined via the usual $\Delta{y}$ versus $v_{\mathrm{drift}}$ method, as already described in Appendix A. Thus, the triangular system of calibration equations for the 8 TPC geometry unkowns is saturated, and the entire TPC system can be calibrated.
After the full alignment correction, drift velocity corection, $t_{0}$ correction, $y_{0}$ correction, the overall systematics of the geometry related calibration parameters of the large TPCs, as estimated via closure tests, are listed in Table 2. The statistical errors are negligible.
@@
Table 2: The table of overall systematics of the TPC geometry parameters, for the large chambers, as estimated via closure tests. The calibration scheme was GRC $\rightarrow$ MTPC-L, followed by MTPC-L $\rightarrow$ VTPC-2, then VTPC-2 $\rightarrow$ MTPC-R and VTPC-2 $\rightarrow$ VTPC-1.

### v2

@@ Appendix B Systematic errors and alignment calibration of the TPCs
which is nothing but the already known Eq.(A.1), re-expressed in the notations of this appendix. Therefore, the remaining corrections $t_{0,\mathrm{corr}}$ and $y_{0}$ can be determined via the usual $\Delta{y}$ versus $v_{\mathrm{drift}}$ method, as already described in Appendix A. Thus, the triangular system of calibration equations for the 8 TPC geometry unknowns is saturated, and the entire TPC system can be calibrated.
After the full alignment correction, drift velocity correction, $t_{0}$ correction, $y_{0}$ correction, the overall systematic uncertainty of the geometry related calibration parameters of the large TPCs, as estimated via closure tests, are listed in Table 2. The statistical uncertainties of the parameters are negligible.
@@
Table 2: The table of overall systematic uncertainties of the TPC geometry parameters, for the large chambers, as estimated via closure tests. The calibration scheme was GRC $\rightarrow$ MTPC-L, followed by MTPC-L $\rightarrow$ VTPC-2, then VTPC-2 $\rightarrow$ MTPC-R and VTPC-2 $\rightarrow$ VTPC-1.